There are 32 vertical fence posts enclosing a rectangular field. The posts are evenly spaced and 2m apart . there are 10 posts along the length of the field . find the area of the field in square meters
A 180 b 252 c 60 d 168 e 308

Answer by checkley77(12844)

L=(10-1)*2=18 METERS
32-2*10=32-20=12 POSTS ALONG THE 2 WIDTHS.
[(12/2)+1]*2=[6+1]2=7*2=14 METERS FOR THE WIDTH.
AREA=L*W
AREA=18*14=252 METERS^2.
ANS. B)